solve the following sysem ofequations.
x+3=5(1)
x=3-3y(2)
What is the solution of the system?

Answer by Maths68(1474)

x+3=5...........(1)
x=3-3y..........(2)
Take (1)
x+3=5...........(1)
x=5-3
x=2
Put the value of x in (2)
x=3-3y..........(2)
2=3-3y
2-3=-3y
-1=-3y
-1/-3=-3y/-3
1/3=y
y=1/3
x,y = 2,1/3
Check
======
Put the value of x and y in (2)
x=3-3y..........(2)
2=3-3(1/3)
2=3-(3/3)
2=3-1
2=2